OpenStudy (anonymous):

Sketch a circle with equation x^2+y^2=1. How many points on the circle have the property that cos s= 0?

OpenStudy (anonymous):

@satellite73 @ParthKohli

OpenStudy (anonymous):

@surjithayer

Parth (parthkohli):

Did you know that the x-coordinate of this circle is cosine?

Parth (parthkohli):

Shall I demonstrate this property?

OpenStudy (anonymous):

no, this is so confusing :/

OpenStudy (anonymous):

yes please

Parth (parthkohli):



Parth (parthkohli):

This is the graph of the circle \(x^2 + y^2 = 1\) (you know how to graph circles, right?)

OpenStudy (anonymous):

I kind of forgot how to graph circles, I haven't done it since geometry, and this is algebra 2.

Parth (parthkohli):

\[(x - h)^2 + (y - k)^2 = r^2\]The above is the general equation of any circle where \((h,k)\) is the centre and \(r\) is the radius.

Parth (parthkohli):

Does that ring a bell?\[x^2 + y^2 = 1\] can be written as\[(x - 0)^2 + (y - 0)^2 = 1^2\]So the centre would be \((0,0)\) and the radius would be \(1\).

OpenStudy (anonymous):

ahh Okay, thanks for refreshing my memory there.

Parth (parthkohli):

The radius is the length of ALL the points falling on the circle from the centre.

Parth (parthkohli):

The points that fall on the circle will be 1 unit away from the centre, as 1 is the radius. So all the points I have illustrated are on the circle, as they are 1 unit away from the centre (0,0).

OpenStudy (anonymous):

Okay so radius =1

Parth (parthkohli):

Now begins the fun part.

Parth (parthkohli):